# Effects of quadratic optomechanical coupling on bipartite entanglements, mechanical ground-state cooling and squeezing in an electro-optomechanical system

### v1

@@ Abstract
We theoretically investigate the steady-state bipartite entanglements, mechanical ground-state cooling, and mechanical quadrature squeezing in a hybrid electro-optomechanical system where a moving membrane is linearly coupled to the microwave field mode of an LC circuit, while it simultaneously interacts both linearly and quadratically with the radiation pressure of a single-mode optical cavity. We show that by choosing a suitable sign and amplitude for the quadratic optomechanical coupling (QOC), one can achieve enhanced and thermally robust stationary bipartite entanglement between the subsystems, improved mechanical ground-state cooling, and $Q$-quadrature squeezing of the mechanical mode beyond the 3-dB limit of squeezing. In particular, we find that in the presence of QOC with positive sign the bipartite optical-microwave entanglement can be increased by about 2.25 times around the temperature of 1mK, and it can be preserved against thermal noise up to the ambient temperature of 90mK. Furthermore, the QOC with positive sign can give rise to the enhancement of the mechanical ground-state cooling by about 1 order of magnitude in the optical and microwave red-detuned regime. We also find that for the positive sign of QOC and near the microwave resonance frequency the squeezing degree of the $Q$-quadrature of the mechanical mode can be amplified up to 7 dB. Such a hybrid electro-optomechanical system can serve as a promising platform to engineer an improved optical- microwave entangled source for microwave sensing as well as quantum information processing.
@@ I Introduction
The optomechanical systems (OMSs) have recently attracted tremendous interest in both theoretical and experimental research in the field of quantum technologies, and have led to a deeper insight of quantum features at macroscopic scale aspelmeyer2014 . These systems, in which the quantized vibration of a mechanical oscillator (MO) is coupled to a quantized electromagnetic (EM) field via a radiation-pressure-type coupling bowen2015 , have been proposed to exploit in an increasingly diverse variety of applications, including cooling of mechanical resonators teufel2011 ; connell2010 ; chan2011 , generation of nonclassical states of the mechanical and optical modes hammerer2014 ; motazedifard2019 ; pontin2016 ; ockeloen2016 ; ockeloen20172 ; pirkkalainen2015 , realization of the optomechanically induced transparency (OMIT) xong2018 ; mikaeili2022 ; agarwal2010 ; weis2010 ; motazedifard2021 ; safavi2011 ; karuza2013 ; kronwald2013 ; motazedifard2022 , generating optomechanical entanglement yang2019 ; palomaki2013 ; paternostro2007 ; dalafi2018 ; barzanjeh2019 ; bemani2019 ; mari2013 ; bemani2017 , ultraprecision quantum sensing and measurements kippenberg2007 ; tsang2010 ; motazedifard2016 ; buchmann2016 ; moler2017 ; liu2019 ; allahverdi2022 ; brunelli2019 ; motazedifard20192 ; ockeloen2018 ; ockeloen20172 ; motazedifardAVS2021 ; ebrahimi2021 ; bemani2021 , quantum illumination or quantum radar barzanjehQIlluminationOMS2015 ; barzanjeh2020 , and detection and interferometry of gravitational waves courty2003 .
@@
It is found that the presence of even a weak QOC can give rise to an enhanced bipartite entanglement between the subsystems. In particular, the steady-state degree of optical-microwave entanglement can be amplified by about 2.25 times around the temperature of 1mK. Moreover, our results reveal that the QOC can also make the enhanced optical-microwave entanglement robust against the thermal noise, such that it survives up to  90 mK. Of particular interest in this respect is that this feature of the system introduces it as an improved quantum radiation source for microwave-quantum illumination target detection. Other interesting findings of the present work concern the role of QOC in enhancing the ground state cooling and quadrature squeezing of the MO. We show that the QOC not only gives rise to a more effective red-sideband cooling of the MO motion to its quantum ground state, but also provides it to start from near sub-Kelvin temperatures without the need of using cryostat. In addition, we find that in the presence of QOC the standard squeezing limit of 3 dB can be beaten for the Q-quadrature of the MO so that it may be squeezed to high degrees (up to 7 dB).
@@ II.1 Mean-field dynamics
Figure 2: (colour online) (a) The mean value of the dimensionless $Q$-quadrature of the MO as a function of the normalized time $\gamma_{m}t$ for $g_{2}=8\times 10^{-3}$ (red), $g_{2}/g_{1}=0$ (grey), and $g_{2}/g_{1}=-8\times 10^{-3}$ (blue). The inset shows the temporal evolution of the $Q$ mean field in a shorter time interval. (b) The mean number of optical cavity photons vs the normalized time $\gamma_{m}t$ for $g_{2}=8\times 10^{-3}$ (red), $g_{2}/g_{1}=0$ (grey), and $g_{2}/g_{1}=-8\times 10^{-3}$ (blue). The optical and microwave cavities are coherently driven by the input fields with powers $\mathcal{P}=30\mu$W and $\mathcal{P}_{w}=30\mu$W, respectively, and the bare detuning parameters have been fixed at $\Delta_{0c}=\Omega_{m}$ and $\Delta_{0w}=-\Omega_{m}$. The other parameters are $\lambda_{c}=810$nm, optical cavity length $\mathcal{L}=1$mm, $m=5$ng, $\Omega_{m}/2\pi=10$MHz, $\mu=0.008$, separation between plates of capacitor $d=100$nm, microwave mode frequency $\omega_{w}/2\pi=10$GHz, decay rate of the optical mode $\kappa=0.01\Omega_{m}$, quality factor of mechanical mode $Q_{m}=8\times 10^{4}$, decay rate of the microwave mode $\kappa_{w}=0.005\Omega_{m}$, and temperature $T=1$mK teufel20112 . The insets show the temporal behaviors of the same mean fields in a shorter time interval.
@@ II.2 Optical multistability
In Fig. 3 we have plotted the steady-state mean number of photons in the optical cavity against the normalized detuning $\Delta_{0c}/\kappa$ in the absence of QOC, i.e., $g_{2}/g_{1}=0$ [Fig. 3] and in its presence with $g_{2}/g_{1}=10^{-3}$ [Fig. 3] and $g_{2}/g_{1}=4\times 10^{-3}$ [Fig. 3]. As is seen from Fig. 3, in the absence of the QOC the system shows an ordinary optical bistability behavior for $\Delta_{0c}>2\kappa$. Here, the optical field is stable along the branches 1 and 3 while it is unstable along the branch 2. The stability conditions can be obtained by using the Routh-Hurwitz criterion dejesus1987 which imposes certain constraints on the system parameters (see the next subsection and Appendix B for the explanation thereof). Figures 3 and 3 exhibit the impact of QOC on the multistability behavior of the system. For the positive values of the QOC parameter we see multistability but when we check the Routh-Hurwitz stability condition only branches 1 and 4 are stable in the system and branches 2, 3, 5 and 6 are unstable for both $g_{2}=10^{-3}$ and $g_{2}=4\times 10^{-3}$. For negative values of the parameter QOC multistability will not occur.
@@ II.3 Stationary quantum fluctuations
Having discussed the mean-field solutions, we now turn our attention to examine the dynamics of quantum fluctuations around the semiclassical steady state. The linearized quantum Langevin equations for the fluctuation operators read as
$\delta\dot{\hat{Q}}=\Omega_{m}\delta\hat{P},$ (12b)
$\qquad-\gamma_{m}\delta P+\sqrt{2\gamma_{m}}\delta P_{in},$ (12c)
where $\tilde{G}=g_{1}+2g_{2}\langle\,Q\rangle_{ss}$ is the QOC-modified optomechanical coupling strength. Here, without loss of generality, we can take $\langle a\rangle_{ss}$ and $\langle a_{w}\rangle_{ss}$ to be real by appropriate choice of the phases of the driving fields $E_{d}$ and $E_{dw}$. By introducing the dimensionless optical-field quadratures fluctuations as $\delta\hat{Q}_{o}=\dfrac{\delta\hat{a}+\delta\hat{a}^{\dagger}}{2}$ and $\delta\hat{P}_{o}=\dfrac{\delta\hat{a}-\delta\hat{a}^{\dagger}}{2i}$, and also the dimensionless microwave-field quadratures fluctuations as $\delta\hat{Q}_{w}=\dfrac{\delta\hat{a}_{w}+\delta\hat{a}_{w}^{\dagger}}{2}$ and $\delta\hat{P}_{w}=\dfrac{\delta\hat{a}_{w}-\delta\hat{a}_{w}^{\dagger}}{2i}$, the set of Eqs. (12) can be expressed in the compact matrix form
@@
where $\delta\hat{u}=(\delta\hat{Q}_{o},\delta\hat{P}_{o},\delta\hat{Q},\delta\hat{P},\delta\hat{Q}_{w},\delta\hat{P}_{w})^{T}$ is the vector of continuous-variable fluctuations and A is the $6\times 6$ drift matrix which is given by
and $\delta n(t)=(\sqrt{2\kappa}\delta\hat{Q}_{in,o},\sqrt{2\kappa}\delta\hat{P}_{in,o},0,\sqrt{2\gamma_{m}}\delta\hat{P}_{in},\sqrt{2\kappa_{w}}\delta\hat{Q}_{in,w}\\ ,\sqrt{2\kappa_{w}}\delta\hat{P}_{in,w})^{T}$ represents the corresponding vector of noises.
The stationary properties of the quantum fluctuations can be examined by considering the steady-state condition governed by Eq. (13). The steady state associated with this equation is reached when the system is stable, which occurs if and only if the real part of all the eigenvalues of the drift matrix A are negative. These stability conditions can be deduced by applying the Routh-Hurwitz criterion which results in six independent inequalities, the explicit expressions of which are given in Appendix B. Because of the linearized dynamics of the quantum fluctuations and since all noises are Gaussian the steady state is a zero-mean Gaussian state which is fully characterized by the 6 ×6 stationary correlation matrix (CM) $V$, with components $V_{i\,j}=\langle\delta u_{i}(\infty)\delta u_{j}(\infty)+\delta u_{j}(\infty)\delta u_{i}(\infty)\rangle/2$. By solving Eq. (13) we obtain
@@
Therefore, the steady-state values of the CM matrix elements are give by
@@
where $\tilde{D}_{kl}(t-t^{\prime})=\langle\delta n_{k}(t)\delta n_{l}(t^{\prime})+\delta n_{l}(t^{\prime})\delta n_{k}(t)\rangle/2$ is the matrix of stationary noise correlation function. Here, $\tilde{D}_{kl}(t-t^{\prime})=D_{kl}\delta(t-t^{\prime})$ in which $D=$ diag$[\kappa,\kappa,0,\gamma_{m}(2n_{m}+1),\kappa_{w}(2n_{w}+1),\kappa_{w}(2n_{w}+1)]$ is a six-dimensional diagonal matrix. The CM in the stationary state is thus obtained as
@@
The effects of QOC on the three kinds of bipartite entanglement are also illustrated in Figs. 5-7, where we have, respectively, plotted $E_{N}^{ow}$, $E_{N}^{om}$, and $E_{N}^{mw}$ versus $\Delta_{0w}/\Omega_{m}$ and $\kappa/\Omega_{m}$ for different values of the ratio $g_{2}/g_{1}$. Figures 5-5 show that in both cases of the absence of the QOC [Fig. 5] and the presence of it [Figs.5 and 5] the logarithmic negativity $E_{N}^{ow}$ is maximized around sideband $\Delta_{0w}=-\Omega_{m}$. However, in the presence of QOC with a positive sign the maximum amount of entanglement between the optical and microwave modes is about twice as large as for the cases without QOC and the QOC with a negative sign. Moreover, as expected, with increasing the optical cavity decay rate $\kappa$, the optical-microwave entanglement decreases drastically. According to Fig. 6, in the absence of QOC there is a weak entanglement between the optical and mechanical modes ($E_{N}^{om}\leq 0.004$) around $\Delta_{0w}=0$. On the Other hand, in the presence of the QOC, whether with positive sign [Fig. 6] or negative sign [Fig. 6], the optical-mechanical entanglement can be strengthened very intensively ($E_{N}^{om}\approx 0.5$). Furthermore, the comparison of Fig. 6 with Figs. 6 and 6 clearly reveals that the presence of the QOC with negative sign leads to a significant robustness of the optical-mechanical entanglement against the optical cavity damping. Concerning to the mechanical-microwave entanglement, we find that the presence of QOC with a positive sign [Fig. 7] reduces while with a negative sign [Fig. 7] does not affect the maximum value of $E_{N}^{mw}$ around sideband $\Delta_{0w}=-\Omega_{m}$ compared to that in its absence [Fig. 7].
@@
As is well known, systems unavoidably suffer from thermal noise, thereby, in general reducing the quantum correlations. The typical investigation in this context is to identify the temperature up to which the entanglement survives which measures its robustness against thermal fluctuations. Here, we are going to explore the influence of the QOC on the thermal robustness of the bipartite entanglements in the system under consideration. For this purpose, we have plotted in Figs. 8-8 the dependence of bipartite logarithmic negativities, $E_{N}^{ow}$, $E_{N}^{om}$, and $E_{N}^{mw}$ on the environment temperature $T$ for different values of the QOC parameter. These figures clearly indicate that the behavior of the entangled subsystem pairs is sensitive to the environment temperature, such that the degree of bipartite entanglement decreases and eventually ceases to exist with increasing environmental temperature. Moreover, the figures show that the effect of QOC manifests as the change in the critical temperature $T_{c}$ which is defined as $T\geqslant T_{c}$, $E_{N}=0$. As is seen from Fig. 8, when the bare microwave detuning is set to $\Delta_{0w}=-\Omega_{m}$ (optimal microwave detuning value that maximizes $E_{N}^{ow}$ , see Fig. 5) the optical-microwave entanglement $E_{N}^{ow}$ in the absence of QOC ($g_{2}=0$) survives until the critical temperature $T_{c}$ reaches about $84m$K. However, in the presence of QOC, either with positive or negative sign, the optical-microwave entanglement exhibits stronger robustness against thermal noise, and the critical temperatures reach as high as about $T_{c}=86m$K and $T_{c}=90m$K for $g_{2}=-4\times 10^{-3}g_{1}$ and $g_{2}=2\times 10^{-3}$, respectively. Figure 8 illustrates the behavior of the optical-mechanical entanglement with respect to temperature where we have set $\Delta_{0w}=-0.4\Omega_{m}$ (optimal microwave detuning value for the optical-mechanical entanglement in the presence of QOC, see Fig. 6). As can be seen, there is no entanglement between the optical and mechanical modes when $g_{2}=0$. Nevertheless, as in the case of the optical-microwave entanglement, the thermal robustness of the optical-mechanical entanglement is enhanced in the presence of QOC. It survives up to $T_{c}~{}110m$K and $T_{c}~{}105m$K for $g_{2}=4\times 10^{-3}g_{1}$ and $g_{2}=-4\times 10^{-3}$, respectively. Lastly, we examine in Fig. 8 the thermal robustness of the mechanical-microwave entanglement $E_{N}^{mw}$ at optimal microwave detuning $\Delta_{0w}=-\Omega_{m}$ (see Fig. 7) for different values of the QOC parameter $g_{2}$. As can be seen, unlike the two former cases, the presence of the QOC decreases the robustness of the entanglement between the mechanical and microwave modes compared to that in its absence. To be more specific, we find that $E_{N}^{mw}$ goes to zero at $T_{c}=96m$K for g2=0 while at $T_{c}=85m$K and $T_{c}=95m$K for $g_{2}=2\times 10^{-3}g_{1}$ and $g_{2}=-4\times 10^{-3}g_{1}$, respectively.
@@
Figure 10: (color online) The effective mean phonon number, $n_{\rm eff}$, vs the dimensionless microwave detuning $\Delta_{0w}/\Omega_{m}$ for different temperatures of the environment. The left panels correspond to $g_{2}=0$, whereas the right ones correspond to $g_{2}=8\times 10^{-3}g_{1}$ in different temperature. All the other parameters are the same as in Fig. 2
Figure 11: (Color online) Plots of effective mean phonon number $n_{\rm eff}$ (on logarithmic scale) and the variances $\langle\delta\hat{Q}^{2}\rangle-3/4$ and $\langle\hat{P}^{2}\rangle-3/4$ versus the normalized detuning $\Delta_{0w}/\Omega_{m}$ for $g_{2}=0$ (panels a-d) and $g_{2}=8\times 10^{-3}g_{1}$ (panels e-h), and for different values of temperature $T=1\rm mk$, $10\rm mK$, $100\rm mK$, $300\rm mK$. Other parameters are the same as in Fig. 2.
@@
where $n_{\rm eff}$ is the mean phonon number of the MO corresponding to an effective temperature $T_{\rm eff}=\hbar\Omega_{m}/[k_{B}\ln(1+1/n_{\rm eff})]$ barzanjeh20112 ; shahidani2014 ; momeni2018 . The ground-state cooling of the MO is achieved whenever $n_{\rm eff}\simeq 0$ or $U\simeq\hbar\Omega_{m}/2$. This can be realized if $\langle\delta\hat{Q}^{2}\rangle=\langle\delta\hat{P}^{2}\rangle\simeq 1/2$ in the steady state. Once the CM $V$ is obtained, one can calculate $n_{\rm eff}$ from its diagonal elements as wang2023

### v2

@@ Abstract
We theoretically investigate the steady-state bipartite entanglements, mechanical ground-state cooling, and mechanical quadrature squeezing in a hybrid electro-optomechanical system where a moving membrane is linearly coupled to the microwave field mode of an LC circuit, while it simultaneously interacts both linearly and quadratically with the radiation pressure of a single-mode optical cavity. We show that by choosing a suitable sign and amplitude for the quadratic optomechanical coupling (QOC), one can achieve enhanced and thermally robust stationary bipartite entanglement between the subsystems, improved mechanical ground-state cooling, and $Q$-quadrature squeezing of the mechanical mode beyond the 3-dB limit of squeezing. In particular, we find that in the presence of QOC with negative sign and in the resolved sideband regime the bipartite optical-mechanical entanglement can be increased by about 2 order of magnitude around the temperature of 1mK, and it can be preserved against thermal noise up to the ambient temperature of 0.1K. Furthermore, the QOC with positive sign can give rise to the enhancement of the mechanical ground-state cooling by about 1 order of magnitude in the optical and microwave red-detuned regime. We also find that for the positive sign of QOC and near the microwave resonance frequency the squeezing degree of the $Q$-quadrature of the mechanical mode can be amplified up to 15 dB. Such a hybrid electro-optomechanical system can serve as a promising platform to engineer an improved entangled source for quantum sensing as well as quantum information processing.
@@ I Introduction
The optomechanical systems (OMSs) have recently attracted tremendous interest in both theoretical and experimental research in the field of quantum technologies, and have led to a deeper insight of quantum features at macroscopic scale aspelmeyer2014 . These systems, in which the quantized vibration of a mechanical oscillator (MO) is coupled to a quantized electromagnetic (EM) field via a radiation-pressure-type coupling bowen2015 , have been proposed to exploit in an increasingly diverse variety of applications, including cooling of mechanical resonators teufel2011 ; connell2010 ; chan2011 , generation of nonclassical states of the mechanical and optical modes hammerer2014 ; pontin2016 ; ockeloen2016 ; ockeloen20172 ; pirkkalainen2015 , realization of the optomechanically induced transparency (OMIT) xong2018 ; mikaeili2022 ; agarwal2010 ; weis2010 ; safavi2011 ; karuza2013 ; kronwald2013 ; motazedifard2022 , generating optomechanical entanglement yang2019 ; palomaki2013 ; paternostro2007 ; dalafi2018 ; barzanjeh2019 ; mari2013 , ultraprecision quantum sensing and measurements kippenberg2007 ; tsang2010 ; buchmann2016 ; moler2017 ; liu2019 ; allahverdi2022 ; brunelli2019 ; ockeloen2018 ; ockeloen20172 ; motazedifardAVS2021 ; ebrahimi2021 ; bemani2021 , quantum illumination or quantum radar barzanjehQIlluminationOMS2015 ; barzanjeh2020 , and detection and interferometry of gravitational waves courty2003 .
@@
It is found that the presence of even a weak QOC can give rise to an enhanced bipartite entanglement between the subsystems. In particular, the steady-state degree of optical-mechanical entanglement can be amplified by about 2 order of magnitude around the temperature of 1mK. Moreover, our results reveal that the QOC can also make the enhanced optical-mechanical entanglement robust against the thermal noise, such that it survives up to 0.1K. Other interesting findings of the present work concern the role of QOC in enhancing the ground state cooling and quadrature squeezing of the MO. We show that the QOC not only gives rise to a more effective red-sideband cooling of the MO motion to its quantum ground state, but also provides it to start from near sub-Kelvin temperatures without the need of using cryostat. In addition, we find that in the presence of QOC the standard squeezing limit of 3 dB can be beaten for the Q-quadrature of the MO so that it may be squeezed to high degrees (up to 15 dB).
@@ II.1 Mean-field dynamics
Figure 2: (colour online) (a) The mean value of the dimensionless $Q$-quadrature of the MO as a function of the normalized time $\gamma_{m}t$ for $g_{2}=8\times 10^{-3}$ (red), $g_{2}/g_{1}=0$ (grey), and $g_{2}/g_{1}=-8\times 10^{-3}$ (blue). (b) The mean number of optical cavity photons vs the normalized time $\gamma_{m}t$ for $g_{2}=8\times 10^{-3}$ (red), $g_{2}/g_{1}=0$ (grey), and $g_{2}/g_{1}=-8\times 10^{-3}$ (blue). The optical and microwave cavities are coherently driven by the input fields with powers $\mathcal{P}=30\mu$W and $\mathcal{P}_{w}=30\mu$W, respectively, and the bare detuning parameters have been fixed at $\Delta_{0c}=\Omega_{m}$ and $\Delta_{0w}=-\Omega_{m}$. The other parameters are $\lambda_{c}=810$nm, optical cavity length $\mathcal{L}=1$mm, $m=5$ng, $\Omega_{m}/2\pi=10$MHz, $\mu=0.008$, separation between plates of capacitor $d=100$nm, microwave mode frequency $\omega_{w}/2\pi=10$GHz, decay rate of the optical mode $\kappa=0.01\Omega_{m}$, quality factor of mechanical mode $Q_{m}=8\times 10^{4}$, decay rate of the microwave mode $\kappa_{w}=0.005\Omega_{m}$, and temperature $T=1$mK teufel20112 . The insets show the temporal behaviors of the same mean fields in a shorter time interval.
@@ II.2 Optical multistability
In Fig. 3 we have plotted the steady-state mean number of photons in the optical cavity against the normalized detuning $\Delta_{0c}/\kappa$ in the absence of QOC, i.e., $g_{2}/g_{1}=0$ [Fig. 3] and in its presence with $g_{2}/g_{1}=10^{-3}$ [Fig. 3] and $g_{2}/g_{1}=4\times 10^{-3}$ [Fig. 3]. As is seen from Fig. 3, in the absence of the QOC the system shows an ordinary optical bistability behavior for $\Delta_{0c}>2\kappa$. Here, the optical field is stable along the branches 1 and 3 while it is unstable along the branch 2. The stability conditions can be obtained by using the Routh-Hurwitz criterion dejesus1987 which imposes certain constraints on the system parameters (see the next subsection and Appendix C for the explanation thereof). Figures 3 and 3 exhibit the impact of QOC on the multistability behavior of the system. For the positive values of the QOC parameter we see multistability but when we check the Routh-Hurwitz stability condition only branches 1 and 4 are stable in the system and branches 2, 3, 5 and 6 are unstable for both $g_{2}=10^{-3}$ and $g_{2}=4\times 10^{-3}$. For negative values of the parameter QOC multistability will not occur.
@@ II.3 Stationary quantum fluctuations
Having discussed the mean-field solutions, we now turn our attention to examine the dynamics of quantum fluctuations around the semiclassical steady state. The linearized quantum Langevin equations for the fluctuation operators can be expressed in the compact matrix form (see Appendix B)
@@
where $\delta{\hat{u}}(t)$ (given by Eq. (30)) is the vector of continuous-variable fluctuations and A (given by Eq. (31)) is the $6\times 6$ drift matrix. Furthermore, $\delta\hat{n}\,(t)$ (given by Eq. (32)) represents the corresponding vector of noises.
The stationary properties of the quantum fluctuations can be examined by considering the steady-state condition governed by Eq. (12). The steady state associated with this equation is reached when the system is stable, which occurs if and only if the real part of all the eigenvalues of the drift matrix A are negative. These stability conditions can be deduced by applying the Routh-Hurwitz criterion which results in six independent inequalities, the explicit expressions of which are given in Appendix C. Because of the linearized dynamics of the quantum fluctuations and since all noises are Gaussian, the steady state is a zero-mean Gaussian state which is fully characterized by the 6 ×6 stationary correlation matrix (CM) V, with components $V_{i\,j}=\langle\delta u_{i}(\infty)\delta u_{j}(\infty)+\delta u_{j}(\infty)\delta u_{i}(\infty)\rangle/2$. By solving Eq. (12) we obtain
@@
Therefore, the steady-state values of the CM matrix elements are given by
@@
where $\tilde{D}_{kl}(t-t^{\prime})=\langle\delta n_{k}(t)\delta n_{l}(t^{\prime})+\delta n_{l}(t^{\prime})\delta n_{k}(t)\rangle/2$ is the matrix of stationary noise correlation function. Here, $\tilde{D}_{kl}(t-t^{\prime})=D_{kl}\delta(t-t^{\prime})$ in which $D=$ diag$[\kappa,\kappa,0,\gamma_{m}(2n_{m}+1),\kappa_{w}(2n_{w}+1),\kappa_{w}(2n_{w}+1)]$ is a six-dimensional diagonal matrix.
The CM in the stationary state is thus obtained as
@@
The effects of QOC on the three kinds of bipartite entanglement are also illustrated in Figs. 5-7, where we have, respectively, plotted $E_{N}^{ow}$, $E_{N}^{om}$, and $E_{N}^{mw}$ versus $\Delta_{0w}/\Omega_{m}$ and $\kappa/\Omega_{m}$ for different values of the ratio $g_{2}/g_{1}$. Figures 5-5 show that in both cases of the absence of the QOC [Fig. 5] and the presence of it [Figs.5 and 5] the logarithmic negativity $E_{N}^{ow}$ is maximized around sideband $\Delta_{0w}=-\Omega_{m}$. However, in the presence of QOC with a positive sign the maximum amount of entanglement between the optical and microwave modes is about 0.04, i.e., about twice as large as for the cases without QOC and the QOC with a negative sign. Moreover, as expected, with increasing the optical cavity decay rate $\kappa$, the optical-microwave entanglement decreases drastically. We point out that the system under consideration may not be optimal for maximizing the entanglement between optical and microwave modes. Nevertheless, the present values of $E_{N}^{ow}$,even if not large, provides the possibility of achieving teleportation fidelities larger than those achievable with only classical means adesso2005 ; mari2008 . It is worth to remark here that it has been theoretically predicted and experimentally reported that higher stationary entanglement between microwave and optical fields (exceeding $10^{-1}$) can be achieved in some platforms, including microwave cavity- optomagnomechanical systems fan2023 , directly coupled electro-optic systems rueda2019 , magnon-based hybrid systems zheng2024 , and optically pulsed superconducting electro-optical devices sahu2023 . According to Fig. 6, in the absence of QOC there is a weak entanglement between the optical and mechanical modes ($E_{N}^{om}\leq 0.004$) around $\Delta_{0w}=0$. On the Other hand, in the presence of the QOC, whether with positive sign [Fig. 6] or negative sign [Fig. 6], the optical-mechanical entanglement can be strengthened very intensively ($E_{N}^{om}\approx 0.5$). More interestingly, the comparison of Fig. 6 with Figs. 6 and 6 clearly reveals that in the presence of the QOC with negative sign the steady-state optomechanical entanglement with considerable robustness against the optical cavity loss is attainable over a relatively wide range of detuning $\Delta_{0w}$ in the resolved sideband regime ($\kappa\,<\Omega_{m}$). This behavior can be understood as follows. According to Eq. (10), in the presence of quadratic coupling the mechanical spring constant is modified from $K=m\Omega_{m}^{2}$ to $K^{\prime}=m\Omega_{m}^{\prime 2}=m\sqrt{\Omega_{m}(\Omega_{m}-2g_{2}|\langle a\rangle_{ss}|^{2})}$ which explicitly depends on the mean value of the intracavity photon number through the parameter $g_{2}$. This so-called optical spring effect can both soften or stiffen the mechanical oscillator depending on the sign of $g_{2}$. On the other hand, the stability of a cavity optomechanical system requires a balance between the radiation-pressure force, given by $F_{r}=\hbar\dfrac{\omega_{c}}{L}|\langle a\rangle_{ss}|^{2}$, and the mechanical restoring force. By using Eqs.(8a), (8b), (9a), (9b), (11a), and (11c), it is found that $F_{r}$ depend complicatedly on the system parameters $\kappa$, $\Omega_{m}$, $g_{2}/g_{1}$, and $\Delta_{0w}$. Since in the presence of QOC with a negative sign the mechanical restoring force increases, the radiation pressure force should also increases accordingly, leading to the enhancement of optomechanical entanglement. Numerical analysis shows that in this situation, the stability range as a function of detuning $\Delta_{0w}$ gets bigger in the resolved sideband regime ($\kappa/\Omega_{m}<1$). In other words, in this regime, the quadratic optomechanical coupling allows for optomechanical entanglement to persist against the cavity loss over a broader range of the detuning $\Delta_{0w}$. Beyond the good cavity limit ($\kappa/\Omega_{m}>1$) this situation is reversed; the larger the cavity damping rate is, the weaker the entanglement observed is. Concerning the mechanical-microwave entanglement, we find that the presence of QOC with a positive sign [Fig. 7] reduces while with a negative sign [Fig. 7] does not affect the maximum value of $E_{N}^{mw}$ around sideband $\Delta_{0w}=-\Omega_{m}$ compared to that in its absence [Fig. 7].
@@
Figure 10: (color online) The effective mean phonon number, $n_{\rm eff}$, vs the dimensionless microwave detuning $\Delta_{0w}/\Omega_{m}$ for different temperatures of the environment. The left panels correspond to $g_{2}=0$, whereas the right ones correspond to $g_{2}=8\times 10^{-3}g_{1}$. All the other parameters are the same as in Fig. 2
As is well known, systems unavoidably suffer from thermal noise, thereby, in general reducing the quantum correlations. The typical investigation in this context is to identify the temperature up to which the entanglement survives which measures its robustness against thermal fluctuations. Here, we are going to explore the influence of the QOC on the thermal robustness of the bipartite entanglements in the system under consideration. For this purpose, we have plotted in Figs. 8-8 the dependence of bipartite logarithmic negativities, $E_{N}^{ow}$, $E_{N}^{om}$, and $E_{N}^{mw}$ on the environment temperature $T$ for different values of the QOC parameter. These figures clearly indicate that the behavior of the entangled subsystem pairs is sensitive to the environment temperature, such that the degree of bipartite entanglement decreases and eventually ceases to exist with increasing environmental temperature. Moreover, the figures show that the effect of QOC manifests as the change in the critical temperature $T_{c}$ which is defined as $T\geqslant T_{c}$, $E_{N}=0$. As is seen from Fig. 8, when the bare microwave detuning is set to $\Delta_{0w}=-\Omega_{m}$ (optimal microwave detuning value that maximizes $E_{N}^{ow}$ , see Fig. 5) the optical-microwave entanglement $E_{N}^{ow}$ in the absence of QOC ($g_{2}=0$) survives until the critical temperature $T_{c}$ reaches about $0.084$K. On the other hand, in the presence of QOC, either with positive or negative sign, the optical-microwave entanglement $E_{N}^{ow}$ at temperatures below the critical temperature is stronger than the case in which $g_{2}=0$. However, as is seen, $E_{N}^{ow}$ disappears at almost the same critical temperature for both cases of $g_{2}=0$ and $g_{2}\neq 0$This means that when the temperature reaches the critical value, the effect of the QOC on $E_{N}^{ow}$ is nearly balanced with that of thermal noise. Figure 8 illustrates the behavior of the optical-mechanical entanglement with respect to temperature where we have set $\Delta_{0w}=-0.4\Omega_{m}$ (optimal microwave detuning value for the optical-mechanical entanglement in the presence of QOC, see Fig. 6). As can be seen, there is no entanglement between the optical and mechanical modes when $g_{2}=0$. Nevertheless, the thermal robustness of the optical-mechanical entanglement is enhanced in the presence of QOC. It survives up to $T_{c}=0.110$K and $T_{c}=0.105$K for $g_{2}=4\times 10^{-3}g_{1}$ and $g_{2}=-4\times 10^{-3}$, respectively. This behavior is different from the one obtained for $E_{N}^{ow}$. That is to say, for the optical- mechanical $E_{N}^{om}$ , the QOC can effectively overcome the effect of thermal noise compared to Fig. 8. Lastly, we examine in Fig. 8 the thermal robustness of the mechanical-microwave entanglement $E_{N}^{mw}$ at optimal microwave detuning $\Delta_{0w}=-\Omega_{m}$ (see Fig. 7) for different values of the QOC parameter $g_{2}$. As can be seen, in contrast to the case of optical-mechanical entanglement, the presence of the QOC decreases the robustness of the entanglement between the mechanical and microwave modes compared to that in its absence.
@@
where $n_{\rm eff}$ is the mean phonon number of the MO corresponding to an effective temperature $T_{\rm eff}=\hbar\Omega_{m}/[k_{B}\ln(1+1/n_{\rm eff})]$ barzanjeh20112 ; shahidani2014 . The ground-state cooling of the MO is achieved whenever $n_{\rm eff}\simeq 0$ or $U\simeq\hbar\Omega_{m}/2$. This can be realized if $\langle\delta\hat{Q}^{2}\rangle=\langle\delta\hat{P}^{2}\rangle\simeq 1/2$ in the steady state. Once the CM $V$ is obtained, one can calculate $n_{\rm eff}$ from its diagonal elements as wang2023